Edelman Electric Systems is considering a project that has the following cash flow and WACC data. What is the project’s MIRR? Note that a project’s projected MIRR can be less than the WACC (and even negative), in which case it will be rejected.


WACC: 10.00%
Year:

0

1

2

3

Cash flows:

–$800

$350

$350

$350

a.

11.82%

b.

9.58%

c.

10.64%

d.

13.14%
Answer #1

Future value of cash flows = Annuity * [(1 + r)n - 1] / r

Future value of cash flows = 350 * [(1 + 0.1)3 - 1] / 0.1

Future value of cash flows = 350 * [1.331 - 1] / 0.1

Future value of cash flows = 350 * 3.31

Future value of cash flows = $1,158.5

MIRR = (Future value / initial investment)1/n - 1

MIRR = (1,158.5 / 800)1/3 - 1

MIRR = (1.4481251/3 - 1

MIRR = 1.1314 - 1

MIRR = 0.1314 or 13.14%